OpenStudy (anonymous):

Type the first equation from the group you selected and identify the slope and the y-intercept of the equation. Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. y = negative three-halves x + 4 here is the equation –3x + 2y = 8 both of them

OpenStudy (anonymous):

i really need help im confused

OpenStudy (calculusfunctions):

What group are you talking about? And what do you mean at the end by both of them?

OpenStudy (anonymous):

thoses are the 2 functions that i have to work with would you like me to put it in a doc so it is easyer to read

OpenStudy (calculusfunctions):

Sure let's see the doc.

OpenStudy (calculusfunctions):

Alright, so your selection is Group 1. Correct?!

OpenStudy (anonymous):

what ever one is better i thought one but what ever one you would like

OpenStudy (calculusfunctions):

It's all the same to me. You make the choice or if you prefer, you can choose one group and I'll select a different group which I'll use as an example to teach you. What do you think? Sound good?

OpenStudy (anonymous):

that seems fine @calculusfunctions

OpenStudy (anonymous):

ok i guess you can teach me group 2 and ill do group one because that is the one that i am assined to by my teacher

OpenStudy (calculusfunctions):

Excellent! Group 2:\[y =x -2\]and\[-2x +3y =-6\] Now the question asks to identify the slope and y-intercept of the first equation:\[y =x -2\]The slope y- intercept form of the equation of a line is given by\[y =mx +b\]where m is the slope of the line and b is the y-intercept of the line. Thus if \[y =x -2\]then are you now able to identify the slope and the y-intercept value?

OpenStudy (anonymous):

i think it might be 2x

OpenStudy (calculusfunctions):

No @mikala1, I said m is the slope> Not mx. Want to try again?

OpenStudy (anonymous):

im not joking when i say i dont understand this i have a vary hard time at math i dont just want the answer like other people i appretiate your helping me and yes plaese

OpenStudy (calculusfunctions):

Alrighty then, no worries! I appreciate your honesty. I'm here for you! It's my job to teach and your job to ask questions and learn. So we have\[y =1x -2\]Now compare that with the general equation\[y =mx +b\]so the slope is m implies that the slope is the coefficient of x (the number in front of the variable x). Hence if\[y =x -2\]which is the same as writing\[y =1x -2\]then the slope must be 1 because in this equation the coefficient of x (the number in front of the variable x) is 1. Let me know when you've had a Eureka! moment.

OpenStudy (anonymous):

oh ok ive got this ok thank you

OpenStudy (calculusfunctions):

Now b is the y-intercept so can you tell me what the y-intercept in that equation would be?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

ok well thank you for ur time i understand alot of it now ik u probeble have other things to do so u can stop helping when ever oyu would like

OpenStudy (calculusfunctions):

Yes but in \[y =x -2\]what is the y-intercept? Sorry my computer lost the connection for a second there.

OpenStudy (calculusfunctions):

Compare the general linear equation with\[y =x -2\]

OpenStudy (anonymous):

x - 2 right since its y = x 2 or 1

OpenStudy (calculusfunctions):

The y-intercept in\[y =mx +b\]is b (the constant or the number independent of the variable x, which means its the term without the x). Thus in \[y =x -2\]the constant term (the term independent of the variable x or the term without the x) is -2 right? Therefore the y-intercept is -2. Do you understand?

OpenStudy (anonymous):

yes i do and u should be a teacher lol ur really good at it

OpenStudy (anonymous):

now on the fist one that you did with me is that the full question and answer because im using that as an example but did we finish it

OpenStudy (calculusfunctions):

Thank you so much? Now let's discuss the next part of the question shall we? Question 3 says: "Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences." Let me ask you if have graph paper in front of you. Do you?

OpenStudy (anonymous):

no she says we have to do it on the computer

OpenStudy (anonymous):

and i said thank you

OpenStudy (calculusfunctions):

Oh I see! Do you know how to?

OpenStudy (anonymous):

no not really i was going to call her for help

OpenStudy (calculusfunctions):

Yes you're very welcome but please there is absolutely no need to say thank you. A teachers greatest reward and thank you is when a student achieves proper understanding and success.

OpenStudy (anonymous):

could you also help me with number 4 for the first one we did please

OpenStudy (calculusfunctions):

The fact that you had a Eureka or a aha moment and now able to answer the questions is what makes me happy. Therefore that is my thanks, which means more than any words you can say.

OpenStudy (calculusfunctions):

Of course! Let me read #4.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

did you leave???

OpenStudy (calculusfunctions):

Again with the thank you. lol Question 4: "Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation." So here the second equation is\[-2x +3y =-6\]Now we are asked to rearrange and express this equation in\[y =mx +b\]form. Do you notice that the left side of the equation has y only and the right side of the equation has the x term and the constant? Good, so then first rearrange\[-2x +3y =-6\]so that there is only the y term (3y) on the left side of the equal sign. To do this all you have to do really is add the opposite of -2x to both sides of the equation. Thus if\[-2x +3y =-6\]then\[3y =2x -6\]Hence we're almost there. Now the reason why it's still incomplete is because we still have 3y on the left side of the equal sign, and we want only. Since 3y means 3 "multiplied by" y, we will divide both sides of\[3y =2x -6\]by 3 (because the inverse operation of multiplication is division, which means division undoes multiplication). Hence if we divide both sides by 3, we obtain\[\frac{ 3y }{ 3 }=\frac{ 2 }{ 3 }x -\frac{ 6 }{ 3 }\]Now simplify to yield\[y =\frac{ 2 }{ 3 }x -2\] Thus the slope of this line is \[\frac{ 2 }{ 3 }\]because that is the number in front of the variable x, and the y-intercept of the line is\[-2\]because that is the term without the x. Do you understand?

OpenStudy (anonymous):

yes that is much simpler than what i thought it would be and another things is im sorry for keeping you but what would be the graphing points for this do you happen to know???

OpenStudy (calculusfunctions):

Sorry, my computer disconnected on me again right in the middle of my response so give me a few minutes because I have to start over again.

OpenStudy (anonymous):

thats ok take ur time no hurry

OpenStudy (calculusfunctions):

They asked us to graph using the the slope-intercept method, right? So i). plot the y-int. For example, in both our equations, the y-int. was -2. Thus first plot the point (0, -2). ii). Now use the slope which is\[\frac{ rise }{ run }\]to help locate a couple or a few more points. In our first equation the slope was 1. This implies that there is a rise of 1 unit and a run of 1 unit. Thus starting from the point of the y-int. simply go 1 unit up and 1 unit to the right to plot the second point. Now from this second point, you may once again, go i unit up and 1 unit right to plot the third plot. You can do this as many times as you wish but just try to stay within the margins of your grid. Note that you may also go 1 unit down and 1 unit left (opposite of 1 up and 1 right) from the point of the y-int. In our second equation, the slope was 2/3. This implies that there is a rise of 2 units up and a run of 3 units to the right. Thus starting from the point of the y-int. go 2 up and 3 right or the opposite of 2 down and 3 left to locate subsequent points. iii). Once you have a few points, simply take a ruler and draw a straight line through passing through the points. Put arrows at the end and label the equation besides the graph. Do you understand?

OpenStudy (anonymous):

thank you so much now only i have to find a ruleer lol

OpenStudy (calculusfunctions):

If that's all for now, I have to go. But before I do, are you absolutely certain, you fully understand the lesson and that you will be fine answering the questions on your own now?

OpenStudy (anonymous):

yes and im sorrry that i took so long to understand it

OpenStudy (calculusfunctions):

You're too polite. There is no need for you to be sorry. If you need my help with anything else involving math then I'll be back later. I'm not sure when bu if you see that I'm on and you need help, just try to catch my attention. Alright? Goodbye for now!

OpenStudy (anonymous):

ok have a great day lol